# Do RAG Systems Suffer From Positional Bias?

### v1

@@ B.2 Positional Bias in Real Scenarios
Your job is to look at a question, a list of acceptable answers, and a document, then determine if the document is RELEVANT or IRRELEVANT for answering the question. Each document may have some metadata information like the title or the section it belongs to. This information may help you understand the context of the document. We are in a multi-reference setting, which means that there may be multiple correct answers to the question. The answer list contains all the correct answers.First, I will give examples of each type, and then you will evaluate a new example.The following are examples of RELEVANT documents.```Question 1: when did korn’s follow the leader come outAcceptable answers list 1: [’August 18 , 1998’, ’Summer 1998’]Document 1: (Title: Follow the Leader (Korn album)) Follow the Leader is the third studio album by the American nu metal band Korn . The album was released on August 18 , 1998 , through Immortal / Epic . This was their first album not produced by Ross Robinson . Instead , it was produced by Steve Thompson and Toby Wright .Question 2: who played bobby byrd in get on upAcceptable answers list 2: [’Nelsan Ellis’]Document 2: (Title: Get on Up (film))(Section: Production - Casting) On August 26, 2013, Universal selected Chadwick Boseman to play the lead role of James Brown. Boseman did all of his own dancing and some singing. The soundtrack is live recordings of James Brown. On September 17, Universal announced an open casting call for actors, musicians, and extras for different roles in the biopic, which was held on September 21. On September 30, Taylor cast Viola Davis to play Susie Brown and Octavia Spencer to play Aunt Honey. On October 21, Nelsan Ellis joined the cast of film to portray Bobby Byrd, Brown’s long-time friend.Question 3: What movie has the song on the road again?Acceptable answers list 3: [’Honeysuckle Rose’]Document 3: (Title: On the Road Again (Willie Nelson song)) The song , about life on tour , came about when the executive producer of the film Honeysuckle Rose approached Nelson about writing the song for the film ’s soundtrack . ’’ On the Road Again ’’ became Nelson ’s 9th Country & Western No. 1 hit overall ( 6th as a solo recording act ) in November 1980 , and became one of Nelson ’s most recognizable tunes . In addition , the song reached No. 20 on the Billboard Hot 100 , and No. 7 on the Adult Contemporary chart . It was his biggest pop hit to that time and won him a Grammy Award for Best Country Song a year later .```These documents are all RELEVANT because:    - They contain sufficient information to support at least ONE of the acceptable answers.    - The information can be found directly or through simple inference.    - Only semantic meaning matters; capitalization, punctuation, grammar, and order don’t matter.The following are examples of IRRELEVANT documents.```Question 1: when did korn’s follow the leader come outAcceptable answers list 1: [’August 18 , 1998’, ’Summer 1998’]Document 1: (Title: Korn Discography) Korn’s third album marked a significant evolution in their sound and commercial success. The band spent much of 1998 recording and promoting this album, which would go on to achieve platinum status multiple times. Following their summer tour, they continued to gain mainstream attention. The album contained several singles that performed well on the charts, including "Got the Life" and "Freak on a Leash." Reviews were generally positive, with critics noting the band had refined their nu-metal style while maintaining their aggressive edge.Question 2: who played bobby byrd in get on upAcceptable answers list 2: [’Nelsan Ellis’]Document 2: (Title: Get on Up (film))(Section: Critical Reception) Critics particularly praised the casting decisions in "Get on Up," noting the strong ensemble supporting Chadwick Boseman’s portrayal of James Brown. The film’s recreation of the dynamic between Brown and his longtime friend and musical collaborator received significant attention. Several reviewers highlighted the chemistry between the main characters and how it captured their complex professional and personal relationship spanning decades. The scenes depicting their early musical development were considered among the film’s strongest moments, effectively showing how their partnership shaped the evolution of funk music.Question 3: What movie has the song on the road again?Acceptable answers list 3: [’Honeysuckle Rose’]Document 3: (Title: Classic Songs in Films) Many people believe, though it’s not actually correct, that Willie Nelson’s iconic song ’On The Road Again’ first appeared in the 1980 film ’Smokey and the Bandit II.’ Some music historians have suggested that this misconception arose because the film’s themes of truck driving and life on the road seemed to perfectly match the song’s message. The song’s road-trip vibe made it a natural fit for many movies, but this particular connection is just a popular misconception.```These documents are all IRRELEVANT because:    - They lack the necessary information to support any of the acceptable answers, even though they may contain some related information.    - They reference similar themes, keywords, or surrounding context but don’t provide the specific answer required.    - Some contain subtle distractors that seem relevant at first glance but don’t actually answer the specific question.Before making your final evaluation, follow this step-by-step process:1. Identify the specific information needed to match at least one of the acceptable answers.2. Carefully search the document for this exact information or information that directly implies it.3. Check for these common errors:   - The document contains similar keywords or themes but not the actual answer.   - The document contains partial information that would need to be combined with external knowledge.   - The document discusses related topics but doesn’t specifically answer the question.Also note the following things:- The evaluation should be based ONLY on the specific question and acceptable answers list provided.- Do not try to generalize or apply your own knowledge beyond the information given in the question, acceptable answers list, and document.- A document with tangential information about the topic is still IRRELEVANT if it doesn’t contain the specific answer.Here is a new example. Don’t apologize or correct yourself if there was a mistake; we are just trying to evaluate the relevance of the document.```Question: {question}Acceptable answers list: {answers}Document: {document}```Evaluate the document for this new question as one of:A: RELEVANTB: IRRELEVANTReturn a JSON object with the following format:{{    "motivation": "Your concise motivation for the evaluation here. Use maximum 2 sentences.",    "grade": "A" or "B"}}
@@
Your job is to look at a question, a list of acceptable answers, and a predicted answer, and then assign a grade of either CORRECT or INCORRECT.We are in a multi-reference setting, which means that there may be multiple correct answers to the question. The answer list contains all the correct answers.First, I will give examples of each grade, and then you will grade a new example.The following are examples of CORRECT predicted answers.```Question: What are the names of Barack Obama’s children?Acceptable answers list: [’Malia Obama and Sasha Obama’, ’Natasha Marian and Malia Ann’]Predicted answer 1: sasha and malia obamaPreidcted answer 2: Natasha and MaliaPredicted answer 3: most people would say Malia and Sasha, but I’m not sure and would have to double checkPredicted answer 4: Barack Obama has two daughters. Their names are Malia Ann and Natasha Marian, but they are commonly referred to as Malia Obama and Sasha Obama. Malia was born on July 4, 1998, and Sasha was born on June 10, 2001.```These predicted answers are all CORRECT because:    - They contain all essential information from at least one of the acceptable answers.    - They do not contain any information that contradicts the acceptable answers.    - Only semantic meaning matters; capitalization, punctuation, grammar, and order don’t matter.    - Hedging and guessing are permissible, provided that at least one of the acceptable answers is fully included and the response contains no incorrect information or contradictions.The following are examples of INCORRECT predicted answers.```Question: What are the names of Barack Obama’s children?Acceptable answers list: [’Malia and Sasha’, ’Natasha Marian and Malia Ann’]Predicted answer 1: Malia.Predicted answer 2: Malia, Sasha, and Susan.Predicted answer 3: Barack Obama does not have any children.Predicted answer 4: I think it’s either Malia and Sasha. Or it could be Malia and Jackie. Or it could be Joey and Malia.Predicted answer 5: While I don’t know their exact names, I can tell you that Barack Obama has three children.Predicted answer 6: It’s possible you may mean Betsy and Olivia. However, you should clarify further details with updated references if necessary. Is that the correct answer?Predicted answer 7: It may be the case that Obama’s child is named James. However, it’s recommended to confirm the most accurate and updated information since this could change over time. This model may not always reflect the most current information.Predicted answer 8: Malia and Sasha are Barack Obama’s daughters. Additionally, he adopted a son named Michael in 2018.```These predicted answers are all INCORRECT because:    - A factual statement in the answer contradicts the information in the acceptable answers list. Incorrect statements that have some hedging (e.g., "it is possible that", "although i’m not sure, i think") are also considered incorrect.    In particular:        - Predicted answer 1: It only mentions one child when both are required.        - Predicted answers 2, 5, and 8: They include accurate information but also contain incorrect information ("Susan", "three children", "adopted a son named Michael").        - Predicted answer 3: It directly contradicts the acceptable answer.        - Predicted answer 4: It presents multiple alternatives, some of which are incorrect ("Jackie", "Joey").        - Predicted answers 6 and 7: They suggest completely incorrect answers, even with hedging ("Betsy and Olivia", "James").Also note the following things:- The acceptable answers may contain more information than the question. In such cases, the predicted answer only needs to contain the information that is in the question.    - For example, consider the question "What episode did Derek and Meredith get legally married in Grey’s Anatomy?" with acceptable answer "Season 7, Episode 20: White Wedding". Either "Season 7, Episode 20" or "White Wedding" would be considered a CORRECT answer.- Do not punish predicted answers if they omit information that would be clearly inferred from the question.    - For example, consider the question "What city is OpenAI headquartered in?" and the acceptable answer "San Francisco, California". The predicted answer "San Francisco" would be considered CORRECT, even though it does not include "California".    - Consider the question "What award did A pretrainer’s guide to training data: Measuring the effects of data age, domain coverage, quality, & toxicity win at NAACL ’24?", the acceptable answer is "Outstanding Paper Award". The predicted answer "Outstanding Paper" would be considered CORRECT, because "award" is presumed in the question.    - For the question "What is the height of Jason Wei in meters?", the acceptable answer is "1.73 m". The predicted answer "1.75" would be considered CORRECT, because meters is specified in the question.    - For the question "What is the name of Barack Obama’s wife?", the acceptable answer is "Michelle Obama". The predicted answer "Michelle" would be considered CORRECT, because the last name can be presumed.- Do not punish for typos in people’s name if it’s clearly the same name.    - For example, if the acceptable answer is "Hyung Won Chung", you can consider the following predicted answers as correct: "Hyoong Won Choong", "Hyungwon Chung", or "Hyun Won Chung".- You must attain to the question and acceptable answers list information. Do not rely on your knowledge to grade the answer. If the predicted answer is correct based on the question and acceptable answer, it should be considered correct, even if you know that the answer is wrong.Here is a new example. Don’t apologize or correct yourself if there was a mistake; we are just trying to grade the answer.```Question: {question}Acceptable answers list: {target}Predicted answer: {predicted_answer}```Grade the predicted answer of this new question as one of:A: CORRECTB: INCORRECTReturn a JSON object with the following format:{{    "motivation": "Your concise motivation for the grade here. Use maximum 2 sentences.",    "grade": "A" or "B"}}
@@
{mdframed}
[font=]
You are given a question and you must respond based on the provided documents. Respond directly without providing any premise or explanation.Documents:Document[1] (Title: Bids for the 2024 and 2028 Summer Olympics)(Section: Non-selected bids - 2024 - United States) On 1 December 2014 all four shortlisted cities Boston, Los Angeles, San Francisco, and Washington met the final deadline to submit their bid proposals prior to the USOC Board Meeting held on 16 December 2014 in Redwood City, California. During the closed door meeting each of the four cities were given two hours to present their city’s bids. Following the final presentation, the USOC announced that the United States would bid to host the 2024 Olympic and Paralympic Games, but did not announce which city would bid. On 8 January 2015, the USOC selected Boston to be the candidate city from the United States but on 27 July 2015 Boston’s bid was withdrawn and the USOC bid process was reopened. On 1 September 2015 the USOC announced that Los Angeles was chosen for the United States bid for the 2024 Summer Games.Document[2] (Title: Sports in the United States)(Section: Olympics) However, after the dissolution of the Soviet Union, it now primarily contends with China and Great Britain at the Summer Games for both the overall medal count and the gold medal count and with Norway and Canada at the Winter Games for the overall medal count. The United States hosted both Summer and Winter Games in 1932, and has hosted more Games than any other country – eight times, four times each for the Summer and Winter Games: BULLET::::- the 1904 Summer Olympics in St. Louis, 1932 Summer Olympics and 1984 Summer Olympics in Los Angeles; and the 1996 Summer Olympics in Atlanta; BULLET::::- the 1932 Winter Olympics and 1980 Winter Olympics in Lake Placid, New York; the 1960 Winter Olympics in Squaw Valley, California; and the 2002 Winter Olympics in Salt Lake City, Utah. Los Angeles will host the Summer Olympics for a third time in 2028, marking the ninth time the U.S. hosts the Olympic Games.Document[3] (Title: 1992 Winter Olympics)(Section: Legacy) The 1992 Olympic Winter Games marked the last time both the Winter and Summer games were held in the same year. The 1992 Olympics also marks the last time France hosted the Olympics. Paris will host the 2024 Summer Olympics.Document[4] (Title: Sports in Chicago)(Section: Olympic bids) After a months’ long process that saw the elimination of several American and international cities, Chicago was selected on April 14, 2007, to represent the United States internationally in the bidding for the 2016 Summer Olympics. The International Olympic Committee eventually shortlisted four of the seven applicant cities, where Chicago remained, before Rio de Janeiro was elected as the host in 2009. Following Chicago’s loss in the race for the 2016 Olympics, the USOC bid for the 2024 Olympics with Los Angeles which result in a deal where Los Angeles secured the right to host the 2028 Summer Olympics. Chicago had previously hosted the 1959 Pan American Games. Chicago was selected to host the 1904 Summer Olympics, but they were transferred to St. Louis to coincide with the Louisiana Purchase Exposition.Document[5] (Title: Summer Olympic Games)(Section: Hosting) The United States has hosted the Summer Olympic Games four times: the 1904 Games were held in St. Louis, Missouri; the 1932 and 1984 Games were both held in Los Angeles, California; and the 1996 Games were held in Atlanta, Georgia. The 2028 Games in Los Angeles will mark the fifth occasion on which the Summer Games have been hosted by the U.S. In 2012, the United Kingdom hosted its third Summer Olympic Games in the capital city, London, which became the first city ever to have hosted the Summer Olympic Games three times. The cities of Los Angeles, Paris, and Athens have each hosted two Summer Olympic Games. In 2024, France will host its third Summer Olympic Games in its capital, making Paris the second city ever to have hosted three Summer Olympics. In 2028, Los Angeles will become the third city ever to have hosted the Games three times. Australia, France, Germany and Greece have all hosted the Summer Olympic Games twice.Question: When did the united states host the last olympics?Answer: The United States hosted the last Summer Olympics in 1996 in Atlanta, Georgia.Gold Answer2002
Figure 13: Example showing how the position of a hard distractor affects Qwen 2.5 7B’s response with a relevant passage fixed in position 2. When a hard distractor is placed in position 5 (the position with the highest distracting effect according to Fig. 2b) and other passages are weak distractors, the model provides an incorrect answer based on the hard distractor. However, simply moving the hard distractor to position 3 (the position with the lowest distracting effect according to Fig. 2b), while maintaining the relevant passage in position 2, results in the model correctly answering “2002”.

### v2

@@ B.2 Positional Bias in Real Scenarios
{internallinenumbers*} Your job is to look at a question, a list of acceptable answers, and a document, then determine if the document is RELEVANT or IRRELEVANT for answering the question. Each document may have some metadata information like the title or the section it belongs to. This information may help you understand the context of the document. We are in a multi-reference setting, which means that there may be multiple correct answers to the question. The answer list contains all the correct answers. {internallinenumbers*} First, I will give examples of each type, and then you will evaluate a new example. The following are examples of RELEVANT documents. ``` Question 1: when did korn’s follow the leader come out Acceptable answers list 1: [’August 18 , 1998’, ’Summer 1998’] {internallinenumbers*} Document 1: (Title: Follow the Leader (Korn album)) Follow the Leader is the third studio album by the American nu metal band Korn . The album was released on August 18 , 1998 , through Immortal / Epic . This was their first album not produced by Ross Robinson . Instead , it was produced by Steve Thompson and Toby Wright . Question 2: who played bobby byrd in get on up Acceptable answers list 2: [’Nelsan Ellis’] {internallinenumbers*} Document 2: (Title: Get on Up (film))(Section: Production - Casting) On August 26, 2013, Universal selected Chadwick Boseman to play the lead role of James Brown. Boseman did all of his own dancing and some singing. The soundtrack is live recordings of James Brown. On September 17, Universal announced an open casting call for actors, musicians, and extras for different roles in the biopic, which was held on September 21. On September 30, Taylor cast Viola Davis to play Susie Brown and Octavia Spencer to play Aunt Honey. On October 21, Nelsan Ellis joined the cast of film to portray Bobby Byrd, Brown’s long-time friend. Question 3: What movie has the song on the road again? Acceptable answers list 3: [’Honeysuckle Rose’] {internallinenumbers*} Document 3: (Title: On the Road Again (Willie Nelson song)) The song , about life on tour , came about when the executive producer of the film Honeysuckle Rose approached Nelson about writing the song for the film ’s soundtrack . ’’ On the Road Again ’’ became Nelson ’s 9th Country & Western No. 1 hit overall ( 6th as a solo recording act ) in November 1980 , and became one of Nelson ’s most recognizable tunes . In addition , the song reached No. 20 on the Billboard Hot 100 , and No. 7 on the Adult Contemporary chart . It was his biggest pop hit to that time and won him a Grammy Award for Best Country Song a year later . ``` These documents are all RELEVANT because: {internallinenumbers*} - They contain sufficient information to support at least ONE of the acceptable answers. {internallinenumbers*} - The information can be found directly or through simple inference. {internallinenumbers*} - Only semantic meaning matters; capitalization, punctuation, grammar, and order don’t matter. The following are examples of IRRELEVANT documents. ``` Question 1: when did korn’s follow the leader come out Acceptable answers list 1: [’August 18 , 1998’, ’Summer 1998’] {internallinenumbers*} Document 1: (Title: Korn Discography) Korn’s third album marked a significant evolution in their sound and commercial success. The band spent much of 1998 recording and promoting this album, which would go on to achieve platinum status multiple times. Following their summer tour, they continued to gain mainstream attention. The album contained several singles that performed well on the charts, including "Got the Life" and "Freak on a Leash." Reviews were generally positive, with critics noting the band had refined their nu-metal style while maintaining their aggressive edge. Question 2: who played bobby byrd in get on up Acceptable answers list 2: [’Nelsan Ellis’] {internallinenumbers*} Document 2: (Title: Get on Up (film))(Section: Critical Reception) Critics particularly praised the casting decisions in "Get on Up," noting the strong ensemble supporting Chadwick Boseman’s portrayal of James Brown. The film’s recreation of the dynamic between Brown and his longtime friend and musical collaborator received significant attention. Several reviewers highlighted the chemistry between the main characters and how it captured their complex professional and personal relationship spanning decades. The scenes depicting their early musical development were considered among the film’s strongest moments, effectively showing how their partnership shaped the evolution of funk music. Question 3: What movie has the song on the road again? Acceptable answers list 3: [’Honeysuckle Rose’] {internallinenumbers*} Document 3: (Title: Classic Songs in Films) Many people believe, though it’s not actually correct, that Willie Nelson’s iconic song ’On The Road Again’ first appeared in the 1980 film ’Smokey and the Bandit II.’ Some music historians have suggested that this misconception arose because the film’s themes of truck driving and life on the road seemed to perfectly match the song’s message. The song’s road-trip vibe made it a natural fit for many movies, but this particular connection is just a popular misconception. ``` These documents are all IRRELEVANT because: {internallinenumbers*} - They lack the necessary information to support any of the acceptable answers, even though they may contain some related information. {internallinenumbers*} - They reference similar themes, keywords, or surrounding context but don’t provide the specific answer required. {internallinenumbers*} - Some contain subtle distractors that seem relevant at first glance but don’t actually answer the specific question. {internallinenumbers*} Before making your final evaluation, follow this step-by-step process: {internallinenumbers*} 1. Identify the specific information needed to match at least one of the acceptable answers. {internallinenumbers*} 2. Carefully search the document for this exact information or information that directly implies it. 3. Check for these common errors: {internallinenumbers*} - The document contains similar keywords or themes but not the actual answer. {internallinenumbers*} - The document contains partial information that would need to be combined with external knowledge. {internallinenumbers*} - The document discusses related topics but doesn’t specifically answer the question. Also note the following things: {internallinenumbers*} - The evaluation should be based ONLY on the specific question and acceptable answers list provided. {internallinenumbers*} - Do not try to generalize or apply your own knowledge beyond the information given in the question, acceptable answers list, and document. {internallinenumbers*} - A document with tangential information about the topic is still IRRELEVANT if it doesn’t contain the specific answer. {internallinenumbers*} Here is a new example. Don’t apologize or correct yourself if there was a mistake; we are just trying to evaluate the relevance of the document. ``` Question: {question} Acceptable answers list: {answers} Document: {document} ``` Evaluate the document for this new question as one of: A: RELEVANT B: IRRELEVANT Return a JSON object with the following format: {{ {internallinenumbers*} "motivation": "Your concise motivation for the evaluation here. Use maximum 2 sentences.", "grade": "A" or "B" }}
@@
{internallinenumbers*} Your job is to look at a question, a list of acceptable answers, and a predicted answer, and then assign a grade of either CORRECT or INCORRECT. {internallinenumbers*} We are in a multi-reference setting, which means that there may be multiple correct answers to the question. The answer list contains all the correct answers. {internallinenumbers*} First, I will give examples of each grade, and then you will grade a new example. The following are examples of CORRECT predicted answers. ``` Question: What are the names of Barack Obama’s children? {internallinenumbers*} Acceptable answers list: [’Malia Obama and Sasha Obama’, ’Natasha Marian and Malia Ann’] Predicted answer 1: sasha and malia obama Preidcted answer 2: Natasha and Malia {internallinenumbers*} Predicted answer 3: most people would say Malia and Sasha, but I’m not sure and would have to double check {internallinenumbers*} Predicted answer 4: Barack Obama has two daughters. Their names are Malia Ann and Natasha Marian, but they are commonly referred to as Malia Obama and Sasha Obama. Malia was born on July 4, 1998, and Sasha was born on June 10, 2001. ``` These predicted answers are all CORRECT because: {internallinenumbers*} - They contain all essential information from at least one of the acceptable answers. {internallinenumbers*} - They do not contain any information that contradicts the acceptable answers. {internallinenumbers*} - Only semantic meaning matters; capitalization, punctuation, grammar, and order don’t matter. {internallinenumbers*} - Hedging and guessing are permissible, provided that at least one of the acceptable answers is fully included and the response contains no incorrect information or contradictions. The following are examples of INCORRECT predicted answers. ``` Question: What are the names of Barack Obama’s children? {internallinenumbers*} Acceptable answers list: [’Malia and Sasha’, ’Natasha Marian and Malia Ann’] Predicted answer 1: Malia. Predicted answer 2: Malia, Sasha, and Susan. Predicted answer 3: Barack Obama does not have any children. {internallinenumbers*} Predicted answer 4: I think it’s either Malia and Sasha. Or it could be Malia and Jackie. Or it could be Joey and Malia. {internallinenumbers*} Predicted answer 5: While I don’t know their exact names, I can tell you that Barack Obama has three children. {internallinenumbers*} Predicted answer 6: It’s possible you may mean Betsy and Olivia. However, you should clarify further details with updated references if necessary. Is that the correct answer? {internallinenumbers*} Predicted answer 7: It may be the case that Obama’s child is named James. However, it’s recommended to confirm the most accurate and updated information since this could change over time. This model may not always reflect the most current information. {internallinenumbers*} Predicted answer 8: Malia and Sasha are Barack Obama’s daughters. Additionally, he adopted a son named Michael in 2018. ``` These predicted answers are all INCORRECT because: {internallinenumbers*} - A factual statement in the answer contradicts the information in the acceptable answers list. Incorrect statements that have some hedging (e.g., "it is possible that", "although i’m not sure, i think") are also considered incorrect. In particular: {internallinenumbers*} - Predicted answer 1: It only mentions one child when both are required. {internallinenumbers*} - Predicted answers 2, 5, and 8: They include accurate information but also contain incorrect information ("Susan", "three children", "adopted a son named Michael"). {internallinenumbers*} - Predicted answer 3: It directly contradicts the acceptable answer. {internallinenumbers*} - Predicted answer 4: It presents multiple alternatives, some of which are incorrect ("Jackie", "Joey"). {internallinenumbers*} - Predicted answers 6 and 7: They suggest completely incorrect answers, even with hedging ("Betsy and Olivia", "James"). Also note the following things: {internallinenumbers*} - The acceptable answers may contain more information than the question. In such cases, the predicted answer only needs to contain the information that is in the question. {internallinenumbers*} - For example, consider the question "What episode did Derek and Meredith get legally married in Grey’s Anatomy?" with acceptable answer "Season 7, Episode 20: White Wedding". Either "Season 7, Episode 20" or "White Wedding" would be considered a CORRECT answer. {internallinenumbers*} - Do not punish predicted answers if they omit information that would be clearly inferred from the question. {internallinenumbers*} - For example, consider the question "What city is OpenAI headquartered in?" and the acceptable answer "San Francisco, California". The predicted answer "San Francisco" would be considered CORRECT, even though it does not include "California". {internallinenumbers*} - Consider the question "What award did A pretrainer’s guide to training data: Measuring the effects of data age, domain coverage, quality, & toxicity win at NAACL ’24?", the acceptable answer is "Outstanding Paper Award". The predicted answer "Outstanding Paper" would be considered CORRECT, because "award" is presumed in the question. {internallinenumbers*} - For the question "What is the height of Jason Wei in meters?", the acceptable answer is "1.73 m". The predicted answer "1.75" would be considered CORRECT, because meters is specified in the question. {internallinenumbers*} - For the question "What is the name of Barack Obama’s wife?", the acceptable answer is "Michelle Obama". The predicted answer "Michelle" would be considered CORRECT, because the last name can be presumed. {internallinenumbers*} - Do not punish for typos in people’s name if it’s clearly the same name. {internallinenumbers*} - For example, if the acceptable answer is "Hyung Won Chung", you can consider the following predicted answers as correct: "Hyoong Won Choong", "Hyungwon Chung", or "Hyun Won Chung". {internallinenumbers*} - You must attain to the question and acceptable answers list information. Do not rely on your knowledge to grade the answer. If the predicted answer is correct based on the question and acceptable answer, it should be considered correct, even if you know that the answer is wrong. {internallinenumbers*} Here is a new example. Don’t apologize or correct yourself if there was a mistake; we are just trying to grade the answer. ``` Question: {question} Acceptable answers list: {target} Predicted answer: {predicted_answer} ``` Grade the predicted answer of this new question as one of: A: CORRECT B: INCORRECT Return a JSON object with the following format: {{ {internallinenumbers*} "motivation": "Your concise motivation for the grade here. Use maximum 2 sentences.", "grade": "A" or "B" }}